OpenStudy (anonymous):

x+square root of x=12

OpenStudy (anonymous):

That's not even the right question Kushash..

OpenStudy (anonymous):

\[x+\sqrt{x}=12\]

OpenStudy (anonymous):

\[\sqrt{x}=12-x=>(\sqrt{x})^2=(12-x)^2=>x=144-24x+x^2\]

OpenStudy (anonymous):

1 subtract x from both side 2 square both side

OpenStudy (anonymous):

my thingy got cut off

OpenStudy (anonymous):

Multiple lines is easier to read.

OpenStudy (anonymous):

do i factor 0=x^2-24x+144

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

what add up to -24 and multiplies to 144?

OpenStudy (anonymous):

kushash.. That is not the right problem. Seriously. Stop spamming.

OpenStudy (anonymous):

kushashwa is half right. x can be 9. he put the answer for root(x).

OpenStudy (anonymous):

i think it should be -25x in the middle

OpenStudy (anonymous):

thats what i thought monkey

OpenStudy (anonymous):

because you still have that other x on the left side

OpenStudy (anonymous):

You can either factor it, or use the quadratic equation.

OpenStudy (anonymous):

right and if you subract that x on both sides you have -24x-x

OpenStudy (anonymous):

in which case it is 0=x^2-25x+144 which makes it more difficult

OpenStudy (anonymous):

x^2-24x-x+144=0

OpenStudy (anonymous):

so then i should use the quadratic equation to solve?

OpenStudy (anonymous):

you can do factoring

OpenStudy (anonymous):

(x-9)(x-16)=0

OpenStudy (anonymous):

x = 9,16

OpenStudy (anonymous):

you guys are awesome thank you

OpenStudy (anonymous):

If it is something you cannot factor easily, using the quadratic formula always works. Even if it's factorable, sometimes you can't figure it out.